# EHRNoteQA: An LLM Benchmark for Real-World Clinical Practice Using Discharge Summaries

### v4

@@ Abstract
Discharge summaries in Electronic Health Records (EHRs) are crucial for clinical decision-making, but their length and complexity make information extraction challenging, especially when dealing with accumulated summaries across multiple patient admissions. Large Language Models (LLMs) show promise in addressing this challenge by efficiently analyzing vast and complex data. Existing benchmarks, however, fall short in properly evaluating LLMs’ capabilities in this context, as they typically focus on single-note information or limited topics, failing to reflect the real-world inquiries required by clinicians. To bridge this gap, we introduce EHRNoteQA, a novel benchmark built on the MIMIC-IV EHR, comprising 962 different QA pairs each linked to distinct patients’ discharge summaries. Every QA pair is initially generated using GPT-4 and then manually reviewed and refined by three clinicians to ensure clinical relevance. EHRNoteQA includes questions that require information across multiple discharge summaries and covers eight diverse topics, mirroring the complexity and diversity of real clinical inquiries. We offer EHRNoteQA in two formats: open-ended and multi-choice question answering, and propose a reliable evaluation method for each. We evaluate 27 LLMs using EHRNoteQA and examine various factors affecting the model performance (e.g., the length and number of discharge summaries). Furthermore, to validate EHRNoteQA as a reliable proxy for expert evaluations in clinical practice, we measure the correlation between the LLM performance on EHRNoteQA, and the LLM performance manually evaluated by clinicians. Results show that LLM performance on EHRNoteQA have higher correlation with clinician-evaluated performance (Spearman($\rho$): 0.78, Kendall($\tau$): 0.62) compared to other benchmarks, demonstrating its practical relevance in evaluating LLMs in clinical settings. EHRNoteQA is publicly available under PhysioNet credential access at https://doi.org/10.13026/acga-ht95, and the code is available at https://github.com/ji-youn-kim/EHRNoteQA.
@@ 1 Introduction
To this end, we present EHRNoteQA, a novel benchmark to evaluate LLMs in real-world clinical scenarios for answering clinicians’ questions regarding patient discharge summaries. EHRNoteQA is built upon the MIMIC-IV EHR [25], and consists of 962 different QA pairs each linked to distinct patients’ discharge summaries. Each QA pair is initially generated using GPT-4 and then carefully reviewed and refined individually by three clinicians to ensure clinical relevance. EHRNoteQA reflects real-world clinician queries for the following reasons: 1) it includes questions that require information across multiple discharge summaries (e.g., “How did the post-discharge medication regimen change between visits x and y?”), 2) the questions span a diverse set of 8 topics (i.e., treatment, assessment, problem, etiology, sign/symptom, vitals, test results, history, instruction, plan). For each question, EHRNoteQA includes a single complete answer along with four distractor answer choices, enabling evaluation of LLMs in both open-ended and multi-choice formats. For both formats, we provide a reliable evaluation method for measuring LLM performance.
@@ 2 Related Works
Previous works have explored patient-specific QA using EHR discharge summaries. Pampari et al. [43] introduced emrQA, the first public clinical note QA dataset generated from expert-annotated question templates and i2b2 annotations [56, 55, 57, 58]. Yue et al. [63] proposed a dataset built on MIMIC-III discharge summaries [24] by extracting answer evidence, followed by generating corresponding questions using a neural question generator model. Additionally, Fan [14] and Moon et al. [37] proposed discharge summary-based QA datasets on why-questions and drug-reason relations, respectively.
@@ 4.3 The Reliability of EHRNoteQA as a Proxy for Clinician Evaluations
Lastly, we reconfirm the validity of our GPT-4-based evaluation method for EHRNoteQA by comparing the correlations between clinician-evaluated LLM scores and EHRNoteQA scores obtained through other evaluation methods. For open-ended questions, when evaluating with traditional methods such as BLEU [44] or ROUGE-L [33], the results show lower correlation compared to our GPT-4 based evaluation scores. For multi-choice questions, many studies employ probability-based scoring. Following this approach, we test two types of probability-based scoring methods on EHRNoteQA using LM-Evaluation Harness [16]: one measuring the probability of the answer index (e.g., A) and another measuring the probability of the correct answer choice text. However, both results show lower correlation compared to our GPT-4 based evaluation scores.
@@ E.1 GPT-4 Evaluation
For evaluating EHRNoteQA on open-ended and multi-choice formats, we utilized the GPT4-turbo-preview (1106) version with default hyperparameters, except for the temperature, which was set to 0. The cost for evaluating the 962 QA data pairs was approximately \$40 per model. The specific prompts for evaluating open-ended and multi-choice formats are shown in Figures 8 and 9. Given the prompt, the model output is evaluated as either correct or incorrect for both open-ended and multi-choice formats (1 for correct, 0 for incorrect).
@@ F.3 Model Performance on Benchmarks
Tables 8 and 9 show the model scores of DiSCQ evaluated by 3 clinicians, along with the scores of different benchmarks including EHRNoteQA. The intra-clinician correlation and the benchmark correlation results in Table 4 are calculated based on the values in Tables 8 and 9.

### v5

@@ Abstract
Discharge summaries in Electronic Health Records (EHRs) are crucial for clinical decision-making, but their length and complexity make information extraction challenging, especially when dealing with accumulated summaries across multiple patient admissions. Large Language Models (LLMs) show promise in addressing this challenge by efficiently analyzing vast and complex data. Existing benchmarks, however, fall short in properly evaluating LLMs’ capabilities in this context, as they typically focus on single-note information or limited topics, failing to reflect the real-world inquiries required by clinicians. To bridge this gap, we introduce EHRNoteQA, a novel benchmark built on the MIMIC-IV EHR, comprising 962 different QA pairs each linked to distinct patients’ discharge summaries. Every QA pair is initially generated using GPT-4 and then manually reviewed and refined by three clinicians to ensure clinical relevance. EHRNoteQA includes questions that require information across multiple discharge summaries and covers ten diverse topics, mirroring the complexity and diversity of real clinical inquiries. We offer EHRNoteQA in two formats: open-ended and multi-choice question answering, and propose a reliable evaluation method for each. We evaluate 27 LLMs using EHRNoteQA and examine various factors affecting the model performance (e.g., the length and number of discharge summaries). Furthermore, to validate EHRNoteQA as a reliable proxy for expert evaluations in clinical practice, we measure the correlation between the LLM performance on EHRNoteQA, and the LLM performance manually evaluated by clinicians. Results show that LLM performance on EHRNoteQA have higher correlation with clinician-evaluated performance (Spearman($\rho$): 0.78, Kendall($\tau$): 0.62) compared to other benchmarks, demonstrating its practical relevance in evaluating LLMs in clinical settings. EHRNoteQA is publicly available under PhysioNet credential access at https://doi.org/10.13026/acga-ht95, and the code is available at https://github.com/ji-youn-kim/EHRNoteQA.
@@ 1 Introduction
To this end, we present EHRNoteQA, a novel benchmark to evaluate LLMs in real-world clinical scenarios for answering clinicians’ questions regarding patient discharge summaries. EHRNoteQA is built upon the MIMIC-IV EHR [26], and consists of 962 different QA pairs each linked to distinct patients’ discharge summaries. Each QA pair is initially generated using GPT-4 and then carefully reviewed and refined individually by three clinicians to ensure clinical relevance. EHRNoteQA reflects real-world clinician queries for the following reasons: 1) it includes questions that require information across multiple discharge summaries (e.g., “How did the post-discharge medication regimen change between visits x and y?”), 2) the questions span a diverse set of 10 topics (i.e., treatment, assessment, problem, etiology, sign/symptom, vitals, test results, history, instruction, plan). For each question, EHRNoteQA includes a single complete answer along with four distractor answer choices, enabling evaluation of LLMs in both open-ended and multi-choice formats. For both formats, we provide a reliable evaluation method for measuring LLM performance.
@@ 2 Related Works
Table 1: Comparison of EHRNoteQA with existing discharge summary QA datasets.
Dataset Questions Documents Patients Answer Format Grounded Documents Topics
Pampari et al. [44] 73,111 303 303 Text Span Single 5
Fan [15] 245 138 138 Text Span Single 1
Yue et al. [65] 1,287 36 36 Text Span Single 5
Moon et al. [38] 96,939 505 505 Text Span Single 1
EHRNoteQA (Ours) 962 1,659 962 Multi-Choice & Open-Ended Multiple 10
@@
Previous works have explored patient-specific QA using EHR discharge summaries (see Table 1 for comparison). Pampari et al. [44] introduced emrQA, the first public clinical note QA dataset generated from expert-annotated question templates and i2b2 annotations [58, 57, 59, 60]. Yue et al. [65] proposed a dataset built on MIMIC-III discharge summaries [25] by extracting answer evidence, followed by generating corresponding questions using a neural question generator model. Additionally, Fan [15] and Moon et al. [38] proposed discharge summary-based QA datasets on why-questions and drug-reason relations, respectively.
@@ 4.3 The Reliability of EHRNoteQA as a Proxy for Clinician Evaluations
Lastly, we reconfirm the validity of our GPT-4-based evaluation method for EHRNoteQA by comparing the correlations between clinician-evaluated LLM scores and EHRNoteQA scores obtained through other evaluation methods. For open-ended questions, when evaluating with traditional methods such as BLEU [45], ROUGE-L [34], Exact Match, and cosine similarity (using SentenceBERT [48] and ClinicalBERT [2]), the results show lower correlation compared to our GPT-4 based evaluation scores. For multi-choice questions, many studies employ probability-based scoring. Following this approach, we test two types of probability-based scoring methods on EHRNoteQA using LM-Evaluation Harness [17]: one measuring the probability of the answer index (e.g., A) and another measuring the probability of the correct answer choice text. However, both results show lower correlation compared to our GPT-4 based evaluation scores.
@@ 5 Discussion
This work was supported by the KAIST-NAVER Hyper-Creative AI Center, the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS-2019-II190075), National Research Foundation of Korea (NRF) grant (NRF-2020H1D3A2A03100945), and Korea Medical Device Development Fund grant (Project Number: 1711138160, KMDF_PR_20200901_0097), funded by the Korea government (MSIT, MOTIE, MOHW, MFDS).
@@ E.1 GPT-4 Evaluation
For evaluating EHRNoteQA on open-ended and multi-choice formats, we utilized the GPT4-turbo-preview (1106) version with default hyperparameters, except for the temperature, which was set to 0. For answer generation (i.e., output) of each model, the temperature was set to 0 for both open-ended and multi-choice formats. The cost for evaluating the 962 QA data pairs was approximately \$40 per model. The specific prompts for evaluating open-ended and multi-choice formats are shown in Figures 8 and 9. Given the prompt, the model output is evaluated as either correct or incorrect for both open-ended and multi-choice formats (1 for correct, 0 for incorrect).
@@ F.3 Model Performance on Benchmarks
Tables 9 and 10 show the model scores of DiSCQ evaluated by 3 clinicians, along with the scores of different benchmarks including EHRNoteQA. The intra-clinician correlation and the benchmark correlation results in Table 5 are calculated based on the values in Tables 9 and 10.